OpenStudy (anonymous):

the sum of three numbers is -4 the second number decreased by the third is equal to the first , the sum of the first and second number is -5, find the numbers

OpenStudy (radar):

Here is an easy start. Let the numbers be x, y, and z Given that x+y=-5 substitute that in the first statement that x + y + z = -4 substituting for x + y -5 + z = -4 z = 1 Now continue on by substituting 1 for z.

OpenStudy (radar):

From this take the statement, "the second number decreased by the third is equal to the first" or y -1 = x, using algebra we can now state: y=x + 1 Substituting the values we now have in the first "given": x + (x + 1) + 1 = -4 2x + 2 = -4 2x = -6 x = -3 Now go 'figger' y as you now know the values of x and z.